Given information for Country X as below:

Details
$ (million)
Consumers consumption =30
Net Investment=20
Government Expenditure=30
Export=20
Import=10
Depreciation=10
Indirect taxes=25
Subsidies=118


1)Calculate Gross Domestic Product (GDP) market price for Country X

2)Calculate Gross Domestic Product (GDP) factor cost for Country X
1
Expert's answer
2020-06-23T10:42:08-0400

C "=" consumption (household) expenditure

I "=" investment expenditure

G "=" government expenditure

x-m "=" net exports that is exports - imports

IT "=" indirect taxes

S "=" subsidies

A) GDP(mp)"=" "C + I + G + (x-m)"

"=30+20+30+(30-20)"

"=90"

B) GDP(fc)"=C + I + G + (x-m) - IT + S"

"=" "30+20+30+(30-20)-25+118"

"=183"